# On the emergence of numerical instabilities in Next Generation Reservoir Computing

### v1

@@ Abstract.
Next Generation Reservoir Computing (NGRC) is a low-cost machine learning method for forecasting chaotic time series from data. However, ensuring the dynamical stability of NGRC models during autonomous prediction remains a challenge. In this work, we uncover a key connection between the numerical conditioning of the NGRC feature matrix — formed by polynomial evaluations on time-delay coordinates — and the long-term NGRC dynamics. Merging tools from numerical linear algebra and ergodic theory of dynamical systems, we systematically study how the feature matrix conditioning varies across hyperparameters. We demonstrate that the NGRC feature matrix tends to be ill-conditioned for short time lags and high-degree polynomials. Ill-conditioning amplifies sensitivity to training data perturbations, which can produce unstable NGRC dynamics. We evaluate the impact of different numerical algorithms (Cholesky, SVD, and LU) for solving the regularized least-squares problem.
Key words and phrases:
Reservoir Computing, Next Generation Reservoir Computing, Forecasting, Structured Matrices, Numerical Linear Algebra, Dynamical Systems
1. Lead Paragraph
@@
Next Generation Reservoir Computing (NGRC) is a reservoir computing (RC) variant characterized by a reduced number of hyperparameters, data efficiency and forecasting quality, and easy interpretability. Despite its success in forecasting tasks, a central challenge remains: ensuring the dynamical stability of the model over long-term forecasts. Recent theoretical work points out a subtle interplay between approximation error and overfitting, yet no analytical method currently exists for selecting hyperparameters that guarantee the NGRC dynamical stability. Based on the premise that numerical instability during training can lead to dynamical instability, we investigate the connection between the choice of hyperparameters and the condition number of the NGRC feature matrix, formed by polynomial evaluations on time-delay coordinates. Notably, the NGRC model features structured matrices, such as Vandermonde-like and Hankel-like matrices, which have been extensively studied in numerical analysis. Merging numerical linear algebra and ergodic theory of dynamical systems, we characterize the conditioning of the feature matrix in terms of the maximum degree of the polynomial basis, the interplay between delay dimension and time lag, and the length of training data. This numerical analysis is a first step toward designing better strategies to select hyperparameters in reservoir computing.
2. Introduction
@@
Forecasting time series is a fundamental problem across scientific disciplines [Cas89, MSW06, McE15], especially when dealing with chaotic systems time series [KS03]. In cases where the governing equations are unknown, the task becomes significantly more challenging. Reservoir Computing (RC) has emerged as a powerful machine learning method for forecasting chaotic dynamical systems directly from data [JH04, LPH+17, LHO18, PLH+17, PHG+18]. A recent promising variant, the Next Generation Reservoir Computing (NGRC) model, recasts the input data as a nonlinear vector autoregressive model [GBGB21]. It requires fewer hyperparameters, shorter training times, and less “warm-up” compared to traditional RC methods [BG22, ZC23]. NGRC stands out for its lower computational cost for forecasting but also other tasks, such as control [KBG24b], reconstruction from partial measurements [GBGB21, RP24], basin reconstruction [GFR22, ZC23] and experimental implementation [KBG24a, CMHR24, WHB+24].
@@
Despite its success in forecasting chaotic time series, a central problem is ensuring the reservoir dynamical stability for short and long-term prediction [Luk12, LHO18, ZC23, GLO25]. Traditionally, in NGRC, training consists of solving a regularized least squares minimization problem (Tikhonov regularization or Ridge regression) to find the readout weight matrix that best predicts one step into the future. The trained model is then run autonomously for forecasting. Although NGRC might be successful during training, NGRC models might be unstable, diverging to infinity during the autonomous testing phase.
@@
Recent theoretical work links reservoir dynamical stability to properties of the original dynamics and the approximation capacity used during training [STY98, DG00, BD23, BD23, GLO25]. By seeing reservoirs as high-dimensional embeddings [HHD20, GHO23], the largest Lyapunov exponent of the reservoir dynamics should not differ much from the original dynamics. However, there is no analytical method to guide hyperparameter selection that guarantees NGRC stability. Developing an analytical method for selecting hyperparameters requires understanding how these choices influence the NGRC dynamics. The mapping from hyperparameters to NGRC dynamics is often nontrivial. For instance, adjusting only the regularization strength [LHO18] can drastically alter the reservoir’s behavior. Moreover, an unstable NGRC model can emerge when the regularizer parameter is not scaled to the length of the training data [ZdSC25].
@@
A promising direction comes from numerical linear algebra: for each hyperparameter configuration, we can examine the condition number of the feature matrix — the matrix that evaluates the input trajectory along the polynomial functions of time-delayed coordinates during training. The condition number quantifies the sensitivity of the least squares solution to perturbations [GVL96, TB97]. When the matrix is ill-conditioned (large condition number), even small changes in the data can produce large readout weights. These large weights are then fed back into the reservoir, potentially driving the model into dynamical instability [ZdSC25]. In this view, numerical instability during training (due to ill-conditioning) can translate into dynamical instability during prediction.
@@
In this paper, we exploit a key structural insight of NGRC: its architecture naturally connects concepts from numerical linear algebra and dynamical systems theory. We show that the NGRC feature matrix possesses well-known structured forms — specifically, a Vandermonde-like matrix, due to the evaluation of polynomial basis functions on time series data, similarly done in polynomial interpolation; and a Hankel-like matrix, due to the use of time-delay coordinates. These matrix classes have been rigorously analyzed in numerical analysis, particularly in terms of their conditioning properties [GVL96, TB97, Hig02, BT17]. By leveraging this structure, we systematically characterize how the conditioning of the feature matrix varies with hyperparameter choices, and we pinpoint the regimes that lead to severe ill-conditioning — a potential precursor to NGRC instability. We use the Lorenz 63 system to demonstrate that the NGRC feature matrix becomes ill-conditioned, particularly for short time lags, high-degree polynomials, and moderate training lengths. Ill-conditioned matrices are a source of numerical instabilities, depending on the condition number and the residual error [GVL96, TB97, Hig02]. Here we test different numerical algorithms to solve the regularized least squares problem of the NGRC training: Cholesky decomposition, SVD, and LU decomposition. Our contribution is not to propose a new method, but to clarify why ill-conditioning arises in NGRC and how it affects long-term prediction stability.
@@
This paper is organized as follows: the preliminaries are presented in Section 3.1. We then introduce the Lorenz system generated by the explicit forward Euler model as a toy model to motivate our problem formulation. Our main results on condition number characterization appear in Section 5. We follow this with numerical testing of different algorithms for solving the regularized least squares problem in Section 6. Finally, we summarize the results and present the discussion in Section 7.
3. Preliminaries
3.1. Notation
[…]
3.2. Learning dynamics
@@
We assume a dynamical system exists $\mathbf{f}:M\to M$ lying on a compact $d-$dimensional metric space $M\subset\mathbb{R}^{d}$. For our purposes, $\mathbf{f}$ is given by discretizing an ordinary differential equation sampled uniformly at every $h$ time step. Instead of knowing $\mathbf{f}$, we only have access to a trajectory $\{\mathbf{x}_{n}\}_{n\geq 0}$ for a given initial condition $\mathbf{x}_{0}$, where $\mathbf{x}_{n}=\mathbf{f}^{n}(\mathbf{x}_{0})$, where $\mathbf{f}^{n}$ denotes the $n$-fold composition of the map $\mathbf{f}$. The problem is to learn the original dynamics $\mathbf{f}$ only from the observed data and be capable of forecasting it forward in time for unseen data.
3.3. Next Generation Reservoir Computing
We focus on using Next Generation Reservoir Computing (NGRC) [GBGB21], which recasts the input data as a nonlinear vector autoregression model [Bol21], akin to a numerical integration scheme. Differently from numerical integration schemes that aim to find the unknown $\{\mathbf{x}_{n}\}_{n\geq 0}$ from a known $\mathbf{f}$, NGRC aims to find an unknown $\mathbf{f}$ from a known $\{\mathbf{x}_{n}\}_{n\geq 0}$.
@@ 3.3. Next Generation Reservoir Computing
(1) $\mathbf{g}_{k,\tau}(\mathbf{x})=\mathbf{vec}\Big{(}\mathbf{x},\mathbf{f}^{\tau}(\mathbf{x}),\dots,\mathbf{f}^{(k-1)\tau}(\mathbf{x})\Big{)},$
that is valid for any point $\mathbf{x}\in M$, and the coordinate in $(\mathbb{R}^{d})^{k}$ is denoted as $\mathbf{X}=\mathbf{g}_{k,\tau}(\mathbf{x})$. Specifically, for the time series $\{\mathbf{x}_{n}\}_{n\geq 0}$, the following holds:
@@
Let $\alpha\in(\mathbb{N}^{d})^{k}$ be the multi-index notation with $|\alpha|=\sum_{j=0}^{k-1}\sum_{i=1}^{d}\alpha_{i,j}$. The set of multivariate monomials in $kd$ variables up to degree $p$ is given by:
@@
(2) $\psi(\mathbf{X}_{n})=\mathbf{vec}\Big{(}\mathbf{X}_{n}^{\alpha}\Big{)}_{|\alpha|\leq p},$
@@
(3) $\mathbf{x}_{n+1}=\mathbf{x}_{n}+\mathbf{W}\psi(\mathbf{X}_{n-(k-1)\tau}),\quad n\geq(k-1)\tau.$
The goal of the training phase is to determine the readout matrix $\mathbf{W}\in\mathbb{R}^{d\times m}$ that best satisfies the above relation at each time step. The available time series $\{\mathbf{x}_{n}\}_{n\geq 0}^{N_{\mathrm{train}}+(k-1)\tau}$ has the number of points necessary to use Eq. 3, where $N_{\mathrm{train}}$ corresponds to the number of training data points and the warm-up $(k-1)\tau$ is the required number of data points to construct the embedded coordinate $\mathbf{X}_{n}$.
For the case of a sufficiently long time series ($N_{\mathrm{train}}>m$), finding $\mathbf{W}$ can be solved by a least squares with a Tikhonov regularization (or Ridge regression). Let us write explicitly the rows of $\mathbf{W}$, such that $\mathbf{W}=(\mathbf{w}_{1}^{\top},\dots,\mathbf{w}_{d}^{\top})$ where $\mathbf{w}_{i}$ is a $m$ dimensional vector. The regression can be made separately for each component of the input trajectory as follows: introduce for $i$-th component
(4) $\mathbf{y}_{i}=\begin{pmatrix}x_{(k-1)\tau+1,i}-x_{(k-1)\tau,i}\\ x_{(k-1)\tau+2,i}-x_{(k-1)\tau+1,i}\\ \vdots\\ x_{(k-1)\tau+N_{\mathrm{train}},i}-x_{(k-1)\tau+N_{\mathrm{train}}-1,i}\end{pmatrix}$
@@
(5) $\Psi=\frac{1}{\sqrt{N_{\mathrm{train}}}}\begin{pmatrix}\psi^{\top}(\mathbf{X}_{0})\\ \psi^{\top}(\mathbf{X}_{1})\\ \vdots\\ \psi^{\top}(\mathbf{X}_{N_{\mathrm{train}}-1})\end{pmatrix}.$
@@
(6) $\mathbf{w}_{i}(\beta)=\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{m}}\Big{\{}\|\mathbf{y}_{i}-\Psi\mathbf{u}\|^{2}+\beta\|\mathbf{u}\|^{2}\Big{\}},$
@@
(7) $\mathbf{w}_{i}(\beta)=(\Psi^{\top}\Psi+\beta\mathbf{1}_{m})^{-1}\Psi^{\top}\mathbf{y}_{i},$
@@
(8) $\mathbf{r}_{n+1}$ $=\mathbf{r}_{n}+\mathbf{W}\psi(\mathbf{R}_{n}),\quad n\geq 0,$
@@
and the autonomous evolution is initialized using the last $(k-1)\tau$ values from the training set to construct the initial embedded coordinate given by the
@@
Figure 1. NGRC model accurately reproduces the Lorenz attractor. The top panel shows the NGRC reconstruction (red) of the Lorenz attractor (blue). The bottom panel shows the NGRC model’s performance over the training and testing phases. The horizontal axis shows time in Lyapunov time $\frac{1}{\Lambda}$ where $\Lambda=0.9056$ is the maximum Lyapunov exponent of the Lorenz system. The parameters are $h=0.01$, delay dimension $k=1$, maximum degree $p=2$, $N_{\mathrm{train}}=500$, $N_{\mathrm{test}}=10000$ (but a smaller time window is shown), and regularizer parameter $\beta=0$.
3.4. Key example: Lorenz system with Explicit forward Euler method
A widely studied chaotic system often used as a benchmark in forecasting tasks for machine learning algorithms is the Lorenz 63 system [Lor63]:
(9) $\begin{split}\dot{x}&=10(y-x)\\ \dot{y}&=x(28-z)-y\\ \dot{z}&=xy-\frac{8}{3}z.\end{split}$

### v2

@@ Abstract
Next Generation Reservoir Computing (NGRC) is a low-cost machine learning method for forecasting chaotic time series from data. Computational efficiency is crucial for scalable reservoir computing, requiring better strategies to reduce training cost. In this work, we uncover a connection between the numerical conditioning of the NGRC feature matrix — formed by polynomial evaluations on time-delay coordinates — and the long-term NGRC dynamics. We show that NGRC can be trained without regularization, reducing computational time. Our contributions are twofold. First, merging tools from numerical linear algebra and ergodic theory of dynamical systems, we systematically study how the feature matrix conditioning varies across hyperparameters. We demonstrate that the NGRC feature matrix tends to be ill-conditioned for short time lags, high-degree polynomials, and short length of training data. Second, we evaluate the impact of different numerical algorithms (Cholesky, singular value decomposition (SVD), and lower-upper (LU) decomposition) for solving the regularized least-squares problem. Our results reveal that SVD-based training achieves accurate forecasts without regularization, being preferable when compared against the other algorithms.
preprint: AIP/123-QED
@@
Next Generation Reservoir Computing (NGRC) is a reservoir computing (RC) variant characterized by a reduced number of hyperparameters, data efficiency and forecasting quality, and easy interpretability. Despite its success in forecasting tasks, a central challenge remains: ensuring the dynamical stability of the model over long-term forecasts in a computationally efficient manner. Recent theoretical work points out a subtle interplay between approximation error and overfitting, yet no analytical method currently exists for selecting hyperparameters that guarantee the NGRC dynamical stability. We investigate the connection between the choice of hyperparameters and the condition number of the NGRC feature matrix, formed by polynomial evaluations on time-delay coordinates. Notably, the NGRC model features structured matrices, such as Vandermonde-like and Hankel-like matrices, which have been extensively studied in numerical analysis. Merging numerical linear algebra and ergodic theory of dynamical systems, we characterize the conditioning of the feature matrix in terms of the maximum degree of the polynomial basis, the interplay between delay dimension and time lag, and the length of training data. This allows us to guide better hyperparameter choices so that an accurate NGRC model can be trained without any regularization, saving computational time. This numerical analysis is a first step toward designing better strategies to select hyperparameters in reservoir computing.
I Introduction
@@
Forecasting time series is a fundamental problem across scientific disciplines [1, 2, 3], especially when dealing with chaotic systems time series [4]. In cases where the governing equations are unknown, the task becomes significantly more challenging. Reservoir Computing (RC) has emerged as a powerful machine learning method for forecasting chaotic dynamical systems directly from data [5, 6, 7, 8, 9]. A recent promising variant, the Next Generation Reservoir Computing (NGRC) model, recasts the input data as a nonlinear vector autoregressive model [10]. It requires fewer hyperparameters, shorter training times, and less “warm-up” compared to traditional RC methods [11, 12]. NGRC stands out for its lower computational cost for forecasting but also other tasks, such as control [13], reconstruction from partial measurements [10, 14], basin reconstruction [15, 12], and experimental implementation [16, 17, 18].
@@
Despite its success in forecasting chaotic time series, a central problem is to ensure computational efficiency for the reservoir dynamical stability for short- and long-term prediction [19, 7, 12, 20]. Traditionally, in NGRC, training consists of fixing a choice of hyperparameter and solving a regularized least squares minimization problem (Tikhonov regularization or ridge regression) to find the readout weight matrix that best predicts one step into the future. The trained model is then run autonomously for forecasting. The successful NGRC model should have a good forecast, not diverging to infinity in the long run. This process must be repeated for each hyperparameter, potentially adding computational cost.
@@
Recent theoretical work links reservoir dynamical stability to properties of the original dynamics and the approximation capacity used during training [21, 22, 23, 23, 20]. By seeing reservoirs as high-dimensional embeddings [24, 25], the largest Lyapunov exponent of the reservoir dynamics should not differ much from the original dynamics. However, there is no analytical method to guide hyperparameter selection that guarantees NGRC stability. Developing an analytical method for selecting hyperparameters requires understanding how these choices influence the NGRC dynamics. The mapping from hyperparameters to NGRC dynamics is often nontrivial. For instance, adjusting only the regularization strength [7] can drastically alter the reservoir’s behavior. Moreover, an unstable NGRC model can emerge when the regularizer parameter is not scaled to the length of the training data [26].
@@
A promising direction comes from numerical linear algebra: for each hyperparameter configuration, we can examine the condition number of the feature matrix — the matrix that evaluates the input trajectory along the polynomial functions of time-delayed coordinates during training. The condition number quantifies the sensitivity of the least squares solution to perturbations [27, 28]. When the matrix is ill-conditioned (large condition number), even small changes in the data can produce large readout weights. These large weights are then fed back into the reservoir, potentially driving the model into dynamical instability [26]. Moreover, ill-conditioning potentially requires regularization, adding computational time to find the regularizer parameter that best forecasts during testing.
@@
In this paper, we exploit a key structural insight of NGRC: its architecture naturally connects concepts from numerical linear algebra and dynamical systems theory. We show that the NGRC feature matrix possesses well-known structured forms — specifically, a Vandermonde-like matrix, due to the evaluation of polynomial basis functions on time series data, similarly done in polynomial interpolation; and a Hankel-like matrix, due to the use of time-delay coordinates. These matrix classes have been rigorously analyzed in numerical analysis, particularly in terms of their conditioning properties [27, 28, 29, 30]. By leveraging this structure, we systematically characterize how the conditioning of the feature matrix varies with hyperparameter choices, and we pinpoint the regimes that lead to severe ill-conditioning. Ill-conditioned matrices are a source of numerical instabilities, depending on the condition number and the residual error [27, 28, 29]. We study chaotic systems to demonstrate that the NGRC feature matrix becomes ill-conditioned, particularly for short time lags, high-degree polynomials, and short training lengths. Using this information, we guide better choices of hyperparameters that allow for NGRC training without any regularization. Furthermore, we test different numerical algorithms to solve the regularized least squares problem of the NGRC training: Cholesky, singular value decomposition (SVD), and lower-upper (LU) decomposition. Our findings show that SVD-based training is more stable and preferable than the other two algorithms.
@@
This paper is organized as follows: the preliminaries are presented in Section˜II.1. We then introduce the Lorenz system generated by the explicit forward Euler model as a toy model in Section˜II.4 to motivate the numerical analysis perspective in Sections˜II.5, II.6 and II.7, and problem formulation in Section˜III. Then, Section˜IV characterizes the condition number of the feature matrix. Subsequently, Section˜V demonstrates two relevant application scenarios where a better choice of hyperparameters allows for an accurate NGRC model without regularization. Finally, we summarize the results and present the discussion in Section˜VI.
II Preliminaries
II.1 Notation
A $d$-dimensional vector is denoted as $\mathbf{u}=(u_{1},\dots,u_{d})$, whereas for a collection of $d$-dimensional vectors, we utilize double indices: $\mathbf{u}_{i}=(u_{i,1},u_{i,2},\dots,u_{i,d})$. The operation $\mathbf{vec}$ concatenates vectors into a single vector, and $\top$ is the transpose. Let $\|\cdot\|$ and $\|\cdot\|_{2}$ be the Euclidean and spectral norm, respectively. The spaces $L^{1}(\mu)$ and $L^{2}(\mu)$ correspond to the space of integrable and square-integrable functions with respect to a probability measure $\mu$, respectively. We denote $\mathbf{1}_{d}$ as the $d\times d$ identity matrix. We utilize Landau’s notation $\mathcal{O}(\varepsilon)$ denoting a function for which there exists a positive constant $K$ such that $0\leq|\mathcal{O}(\varepsilon)|\leq K\varepsilon$ for the limit $\varepsilon\to 0$ or $\varepsilon\to\infty$, being clear by the context.
II.2 Learning dynamics
@@
We assume a dynamical system exists $\mathbf{f}:M\to M$ lying on a compact $d$-dimensional metric space $M\subset\mathbb{R}^{d}$. For our purposes, $\mathbf{f}$ is given by discretizing an ordinary differential equation sampled uniformly at every $h$ time step. Instead of knowing $\mathbf{f}$, we only have access to a trajectory $\{\mathbf{x}_{n}\}_{n\geq 0}$ for a given initial condition $\mathbf{x}_{0}$, where $\mathbf{x}_{n}=\mathbf{f}^{n}(\mathbf{x}_{0})$, where $\mathbf{f}^{n}$ denotes the $n$-fold composition of the map $\mathbf{f}$. The observed time series is normalized coordinate-wise by applying the transformation $\mathbf{x}_{n,i}\mapsto\frac{\mathbf{x}_{n,i}}{\max_{n}\{\mathbf{x}_{n,i}\}}$ for each coordinate $i=1,\dots,d$. The problem is to forecast the original dynamics forward in time only from the observed data.
II.3 Next Generation Reservoir Computing
We focus on using Next Generation Reservoir Computing (NGRC) [10], which recasts the input data as a nonlinear vector autoregression model, akin to a numerical integration scheme. This particular architecture originated from the mathematical equivalence between a traditional RC and the NGRC: there exists an NVAR that performs equally as well as an optimized RC [31]. The key observation is that RC can be recast using a linear activation function and nonlinear readouts, akin to a Volterra series, which have been proven to be a universal approximator [32]. Although this equivalence only occurs for infinite memory, in practice, truncated finite memory suffices, speeding up training and making it the main advantage of the NGRC. Interestingly, NGRC has also been shown to be a special case of polynomial kernel regression using time-delayed coordinates [33], showing its relevance inside the broader context of machine learning approaches to dynamical systems.
@@ II.3 Next Generation Reservoir Computing
$\mathbf{g}_{k,\tau}(\mathbf{x})=\mathbf{vec}\Big(\mathbf{x},\mathbf{f}^{\tau}(\mathbf{x}),\dots,\mathbf{f}^{(k-1)\tau}(\mathbf{x})\Big),$ (1)
which is valid for any point $\mathbf{x}\in M$, and the coordinate in $(\mathbb{R}^{d})^{k}$ is denoted as $\mathbf{X}=\mathbf{g}_{k,\tau}(\mathbf{x})$. Specifically, for the time series $\{\mathbf{x}_{n}\}_{n\geq 0}$, the following holds:
@@
Let $\alpha\in(\mathbb{N}^{d})^{k}$ be the multi-index vector with $|\alpha|=\sum_{j=0}^{k-1}\sum_{i=1}^{d}\alpha_{i,j}$. The set of multivariate monomials in $kd$ variables up to degree $p$ is given by:
@@
$\psi(\mathbf{X}_{n})=\mathbf{vec}\Big(\mathbf{X}_{n}^{\alpha}\Big)_{|\alpha|\leq p},$ (2)
@@
$\mathbf{x}_{n+1}=\mathbf{x}_{n}+\mathbf{W}\psi(\mathbf{X}_{n-(k-1)\tau}),\quad n\geq(k-1)\tau.$ (3)
The goal of the training phase is to determine the readout matrix $\mathbf{W}\in\mathbb{R}^{d\times m}$ that best satisfies the above relation at each time step. The available time series $\{\mathbf{x}_{n}\}_{n\geq 0}^{N_{\mathrm{train}}+(k-1)\tau}$ has the number of points necessary to use Eq.˜3, where $N_{\mathrm{train}}$ corresponds to the number of training data points and the warm-up $(k-1)\tau$ is the required number of data points to construct the embedded coordinate $\mathbf{X}_{n}$.
For the case of a sufficiently long time series ($N_{\mathrm{train}}>m$), finding $\mathbf{W}$ can be solved by a least squares with a Tikhonov regularization (or ridge regression). Let us write explicitly the rows of $\mathbf{W}$, such that $\mathbf{W}=(\mathbf{w}_{1}^{\top},\dots,\mathbf{w}_{d}^{\top})$ where $\mathbf{w}_{i}$ is a $m$ dimensional vector. The regression can be made separately for each coordinate of the input trajectory as follows: introduce for $i$-th coordinate
$\mathbf{y}_{i}=\begin{pmatrix}x_{(k-1)\tau+1,i}-x_{(k-1)\tau,i}\\ x_{(k-1)\tau+2,i}-x_{(k-1)\tau+1,i}\\ \vdots\\ x_{(k-1)\tau+N_{\mathrm{train}},i}-x_{(k-1)\tau+N_{\mathrm{train}}-1,i}\end{pmatrix}$ (4)
@@
$\Psi=\frac{1}{\sqrt{N_{\mathrm{train}}}}\begin{pmatrix}\psi^{\top}(\mathbf{X}_{0})\\ \psi^{\top}(\mathbf{X}_{1})\\ \vdots\\ \psi^{\top}(\mathbf{X}_{N_{\mathrm{train}}-1})\end{pmatrix}$. (5)
@@
$\mathbf{w}_{i}(\beta)=\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{m}}\Big\{\|\mathbf{y}_{i}-\Psi\mathbf{u}\|^{2}+\beta\|\mathbf{u}\|^{2}\Big\},$ (6)
@@
$\mathbf{w}_{i}(\beta)=(\Psi^{\top}\Psi+\beta\mathbf{1}_{m})^{-1}\Psi^{\top}\mathbf{y}_{i},$ (7)
@@
$\mathbf{r}_{n+1}$ $=\mathbf{r}_{n}+\frac{1}{\sqrt{N_{\mathrm{train}}}}\mathbf{W}\psi(\mathbf{R}_{n}),\quad n\geq 0,$ (8)
@@
and the autonomous evolution is initialized using the last $(k-1)\tau$ values from the training set to construct the initial embedded coordinate given by
@@
Figure 1: NGRC model accurately reproduces the Lorenz attractor. The top panel shows the NGRC reconstruction (red) of the Lorenz attractor (blue). The bottom panel shows the NGRC model’s performance over the training and testing phases. The horizontal axis shows time in Lyapunov time $\frac{1}{\Lambda}$ where $\Lambda=0.9056$ is the maximum Lyapunov exponent of the Lorenz system. The parameters are $h=0.01$, delay dimension $k=1$, maximum degree $p=2$, time lag $\tau=1$, $N_{\mathrm{train}}=500$, $N_{\mathrm{test}}=10000$ (but a smaller time window is shown), and regularizer parameter $\beta=0$.
II.4 Key example: Lorenz system with explicit forward Euler method
A widely studied chaotic system often used as a benchmark in forecasting tasks for machine learning algorithms is the Lorenz 63 system [34]:
$\begin{split}\dot{x}&=10(y-x)\\ \dot{y}&=x(28-z)-y\\ \dot{z}&=xy-\frac{8}{3}z.\end{split}$ (9)